Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS September/October 2010  Not logged in ELOG logo
Message ID: 77     Entry time: Mon Oct 4 14:15:22 2010
Author: Manfred Valentan 
Project: Pt 
Subject: run 
Run Number: run040 --> unusable --> no bias 
Events: 25600 
StartTime: 04.10.2010 14:?:? 
EndTime: 04.10.2010 14:32:14 

Pt modules, angle scan

h01-Pt and h02-Pt
currents: h01=26.0A h02=13.4A
HV=200V beam on
angle: 47 deg

telescope run 20146

bias voltage has been ramped down during run039 --> needs to be repeated.

ELOG V3.1.4-966e3dd